# Leveraging Interpolation Models and Error Bounds for Verifiable Scientific Machine Learning

### v1

@@ Abstract
Keywords: interpolation; scientific machine learning; verification and validation; uncertainty quantification
@@ 1 Introduction
To better understand the performance of interpolants on large, complex data sets, we turn to the classical techniques for multivariate interpolation via nonparametric models. One of the main advantages to restricting our study to interpolation as opposed to including regression, is that many interpolation methods have strict and easily-interpretable error bounds and well-known convergence rates [13]. In the context of scientific machine learning, these error bounds can be leveraged to verify the correctness of a model and its applicability to a given problem. Additionally, the direct dependence of these models and their bounds on the training data or simple basis functions, allows for predictions to be interpreted, particularly when working with small- to medium-sized data sets.
@@ 2 Interpolation Methods and Error Bounds
In this section we discuss common methods for interpolating an unknown blackbox function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ given arbitrary training data inputs ${\cal D}=\{x_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$ and paired outputs $f\left(\cal D\right)=\{y_{i}\}_{i=1}^{n}\subset\mathbb{R}$ at an arbitrary “query” or “test” point $q\in{\cal X}\subset\mathbb{R}^{d}$ (but $q\not\in{\cal D}$). Specifically, we have chosen common methods from the literature that are well-defined for large values of $n$ and $d$, and do not rely upon any structured arrangement of ${\cal D}$. All of our analysis is also based upon scattered training data ${\cal D}$, which is the least restrictive case.
@@ 2.1.2 Advantages and challenges of Delaunay interpolants
A challenge when employing Delaunay interpolants is a potentially expensive computational cost at inference time. Many common approaches for computing Delaunay triangulations are impractical in even moderate dimensions, due to the exponential size of $DT({\cal D})$ with growing $d$ [35]. However, it is possible to identify the single simplex needed to calculate ${\hat{f}}_{DT}(q)$ for any given $q$ at inference time, for the cost of solving a linear program [36, 37]. This approach is scalable enough for usage with moderately-sized training sets, but still considerably more expensive than making a single prediction with many comparable neural network models, such as MLPs. Additionally, when implementing such an approach, one must be careful of geometric degeneracy, which results in a degenerate linear program and could cause failures in certain solvers. These degeneracies occur with probability zero in theory, but are extremely common in real-world data sets. For instance, if any subset of a training set is grid-aligned, the geometry of the data would be considered degenerate.
@@ 2.2 RBF Interpolants and Thin-Plate Splines
The first step in fitting an RBF interpolant is (typically) fitting a “polynomial tail” $t(q)$, computed via generalized linear regression with polynomial basis functions. In this paper, we will use either a least-squares approximation for a linear tail, resulting in $t(q)=t_{1}(q):=\alpha^{\top}q+\beta$ for some $\alpha,\beta$, or a constant tail given by $t(q)=t_{0}(q):=\sum_{i=1}^{n}f(x^{(i)})/n$.
@@ 2.2.1 Error bounds for RBFs
Error analysis for RBF interpolation is tricky because the error bounds depend on globally-dependent quantities derived from $f$, which would rarely be known in practice. We focus on a bound provided by Powell [40] and explained in [41, Theorem 1] as follows. Define
@@
Sticking with a mere Lipschitiz continuity assumption for $f$, we are not aware of a computable estimate for $C$ or $\|f\|_{\Phi}$. However, it is useful to note the scaling behavior of (10) with respect to $h$. Note that $h$ is a measure of the density of ${\cal D}$ in ${\cal X}$. If ${\cal D}$ becomes dense enough, the estimate is dominated by the $\log(h^{-1})$ term, which blows up as $h\rightarrow 0$, in that large $h$ values indicate less density, somehwere in $\mathcal{X}$. The blow up is a feature, not a bug, of the estimate, as it reflects the numerical ill-conditioning of the linear system (8) as the training data clusters.
@@ 2.2.2 Advantages and challenges of RBFs
When the size of the data set $n$ is small, RBF interpolants are relatively cheap to fit. Additionally, many choices of basis functions—including the thin-plate spline functions that we will describe in the next section—there are often no hyperparameters. Evaluation of the interpolant at query-time is also quick with $O(nd)$ complexity. It is thus no surprise that RBF methods are a popular off-the-shelf solution to a wide variety of interpolation problems.
@@
Compared to the other interpolants that we consider, one critical challenge for a generic RBF interpolant is that the error bound cannot be easily calculated or approximated. One critical unknown in (10 is the value of the constants $C$ and $\|f\|_{\Phi}$. In principle, these values can be determined independently of the data set ${\cal D}$, however in practice they cannot be known for many real-world applications. We will see a special case in Section 2.3.1 where an error bound can be derived by statistical means, but this approach is kernel-dependent.
@@ 2.3.1 Approximate error bounds for GPs
Following the analysis in [16, Ch. 2], the variance of the posterior distribution for ${\hat{f}}_{GP}$ at $q$ is given by
$p\left({\hat{f}}_{GP}(q)~{}|~{}{\cal D}\right)\sim{\cal N}\left(f;\mathbb{E}\left[{\hat{f}}_{GP}\right](q),K(q,x)\right),$
@@
$K(q,x)=\tau_{f}\left(e^{-\|q-x\|_{2}^{2}/\tau}-\left(\phi^{(0)}(q),\ldots,\phi^{(1)}(q)\right)A_{GP}^{-1}\left[\begin{array}[]{c}\phi^{(0)}(x)\\ \vdots\\ \phi^{(1)}(x)\end{array}\right]\right),$
where $\tau_{f}$ is the true variance in response values for $f$ and $A_{GP}$ is given by the RBF interpolation matrix from (8) with $\Phi=GP$. For an estimate of the variance at $q$, which can be used to evaluate the expected variance in $|f(q)-{\hat{f}}_{GP}(q)|$, this reduces to
$K(q,q)=\tau_{f}\left(1-\left\|\left(\phi^{(0)}(q),\ldots,\phi^{(1)}(q)\right)\right\|_{A^{-1}_{GP}}^{2}\right)$ (12)
where $\|\cdot\|_{A^{-1}_{GP}}$ is the 2-norm rescaled by $A^{-1}_{GP}$, which is also a norm since $A^{-1}_{GP}$ is symmetric positive definite.
Note that this interpretation of GPs relies on additional assumptions about the smoothness of $f$ that may not hold true in practice. For many real-world $f$, it is incorrect to assume that response values are Gaussian distributions and correlated with other response values based on the exponential of the inverse squared distance in the input space. Further, if $f$ is not smooth enough, the error approximation by the above interpretation yields poor estimates. Nevertheless, the uncertainty estimate is widely used and is considered a useful heuristic in many applications.
@@ 2.3.3 Our methods for fitting GP interpolants
Instead, we solve a small well-defined optimization problem based on the statistical literature to select a maximum likelihood estimate of $\tau$ via scikit-learn [46]. Specifically, the log-marginal likelihood for observing the response values $\{y_{i}\}_{i=1}^{n}$ given the data $\{x_{i}\}_{i=1}^{n}$ is
$\log p(y_{i}|x_{i})=-\frac{1}{2}\sum_{i=1}^{n}y_{i}c_{i}-\sum_{i=1}^{n}\log L_{ii}-\frac{n}{2}\log 2\pi,$
where $L$ is a Cholesky factor of $A_{GP}$ [47]. In scikit-learn, this marginal likelihood function is maximized over possible values of $\tau$ using a multi-start L-BFGS solver. While this is still technically a small hyperparameter optimization problem, the problem is relatively inexpensive to solve and extremely reproducable since we are tuning a single variable using second-order gradient-based methods.
@@ 2.4 Artificial Neural Network and Multilayer Perceptron Models
where $u(\cdot)$ is a nonlinear activation function, $W^{(i)}$ is a tensor who’s dimensions are compatible with $x^{(i)}$, and $b^{(i)}$ is a tensor whose dimension match $W^{(i)}x^{(i)}$.
@@ 3.1 Dataset generation and experimental setup
We have carried out numerous experiments exploring the interpolant errors and bounds over this parameter space and report the most relevant findings below.
@@ 3.2 Effect of function variation ($\omega$) over $n$ and $d$
In Figure 2, we show the results for fixing $n=8192$, with additional conventions as in the previous figure. For the Delaunay and MLP results, notice that the relative positions of the data are mostly stable for $d\geq 5$. We infer that the $d=5$ case from Figure 1a is representative of the relative performance of these interpolants for higher dimensions (at least up to 20). For $d<5$, we see smaller errors in the interpolation cases, due to the relative density of our fixed $n$ in low dimensions. As $d$ grows, the error effectively plateaus while the bound increases, reflecting the greater uncertainty inherent in the higher dimensional settings.
@@ 3.3 Effect of skew ($\alpha$) and the importance of scaling
3.4 Prevalence of extrapolation regimes as $d$ increases
@@
In Figure 5, we show the results of the following experiment. For $n=2^{8+2i}$, $i=0,1,2,3$ (the same values used in Figures 13, and 4) we create a data set using a Sobol sequence of $n$ points in $[-1,1]^{d}$ for $d=2,5,8,11$. We set $\alpha=0$ (no skew). The value of $\omega$ is irrelevant as response values are not used in this study. For each $d$, we draw $2^{10}$ points uniformly from $[-1,1]^{d}$, which we call the “test points.” For each test point $q$, we compute $||q||$ and determine whether $q$ lies inside or outside the convex hull of the data set, corresponding to an interpolation or extrapolation query, respectively. Finally, we plot a histogram of the $||q||$ values, using 20 bins from 0 to $\sqrt{d}$ and color each bar according to the proportion of interpolation (blue) vs extrapolation (red) queries.
@@
Additionally, observe that as dimension increases, the distribution of $||q||$ values becomes more sharply clustered in the interior of the range of possible values (note $||q||\in\left[0,\sqrt{d}\right]$ for $q\in[-1,1]^{d}$). This phenomenon, known as concentration of volume, is sometimes overlooked in the data science community although it is well understood from the mathematical setting; see, for instance, [62, Theorem 3.2] and the references that follow. The concentration in $||q||$ values is due in part to the uniform sampling strategy we selected to draw the test points, however, we observed very similar results when using Latin hypercube or Sobol sequence sampling instead. Weighted sampling strategies could help distribute $||q||$ values more evently, however, whether that is beneficial or advisable depends highly on the application context.
@@ 4 Case Study: Predicting Lift-Drag Ratios from Airfoil Images
Our training data for this task is obtained from the UIUC Airfoil Coordinates Database[63]. This is a collection of coordinates for approximately 1,600 airfoils and their corresponding lift-to-drag ratios. These data correspond to a wide range of applications from low Reynolds number airfoils for UAVs and model aircraft to jet transports and wind turbines. Therefore, rapid predictions for the quantity of interest for this dataset corresponds to potential savings in predictive modeling across several different domains.
The input data is represented as a series of pixels having value 0 for points on or within the airfoil and value 1 for points in the free-stream surrounding the airfoil. The output data is a scalar variable for the lift-to-drag ratio. We leverage a publicly available version of the data  [^3] https://github.com/ziliHarvey/CNN-for-Airfoil, thus demonstrating a purely data-driven approach for predicting the outputs. Specifically, this repository deploys a deep learning convolutional neural network to ingest two-dimensional representations of the airfoil coordinates to directly predict the lift-to-drag ratio.
One of the key challenges of this problem from an interpolation perspective is the high dimensionality of the problem when considered in the pixel space. Modern deep learning approaches for image classification would embed these inputs into a lower-dimensional latent space while exploiting the spatial and scale-invariant structures of the image via either convolutional layers or vision transformers. Since our problem involves regression, not classification, we will employ a combination of embedding layers and ReLU MLP-esque regression layers to implement a scientific machine learning approach. In order to also leverage interpolation techniques, we have broken up the training of these layers (denoted $f_{R}$ and $f_{MLP}$ in Section 2.4).
In our approach, we first utilize a convolutional autoencoder to compress the two-dimensional input data to a low-dimensional latent space (i.e., $f_{R}$ in the language of Section 2.4). The latent representations of our airfoils are then used for approximating the lift-to-drag ratio using the various interpolation methods. Our convolutional autoencoder is trained on 80% of the total data set, using a mean-squared error loss function. Latent space representations are obtained using the trained encoder, which are applied on the raw input data. Specific details of the autoencoder architecture can be observed in our associated code. The choice of latent space dimension, validation, and testing are described in subsequent sections. The remaining points are reserved for a validation and testing set, described in the next subsections.
Note that for all interpolation techniques and the ReLU MLP, we have used the same implementations described in Section 3 with input and output rescaling.
4.1 Validation
To validate our latent space encoding and models, we interpolate the 5,553 training points used to train the convolutional autoencoder, and measure the resulting errors for 616 validation points. In order to visualize the convergence of our methods when adding training data, we measure the relative mean absolute error (MAE) of each interpolation method and the ReLU MLP on the validation set using various percentages of the $5,553$ training points. In addition, we have measured the average error bounds across all training sizes. In order to visualize the effectiveness of the error bounds in predicting actual error, we also display a scatter plot of the predicted error bounds vs. observed error for all validation data when using the full training set. The resulting relative errors are shown for a 2-dimensional latent space encoding in Figure 6, a 4-dimensional latent space in Figure 7, a 6-dimensional latent space in Figure 8, and a 8-dimensional latent space in Figure 9.

### v2

@@ 1 Introduction
A sample application—which we will consider via a detailed case study later in the paper—is the prediction of the lift-to-drag ratio for an airfoil given only an image of its cross-section. Training data in this context is a collection of lift-to-drag ratios for some airfoil designs, determined via computationally expensive numerical simulations of the Navier-Stokes equations. Encoding the space of airfoil cross section images into a latent space $\mathbb{R}^{d}$ for some small $d$ is now standard machine learning practice via autoencoder methods. We will show that analyzing the performance of standard interpolation techniques on the encoded images—i.e., from $\mathbb{R}^{d}$ to $\mathbb{R}^{1}$—reveals unique information about the validity of the trained autoencoder and enables simple visual interpretations of its failures.
One of the main advantages to restricting our study to interpolation as opposed to including regression, is that many interpolation methods have strict and easily-interpretable error bounds and well-known convergence rates [13]. In the context of scientific machine learning, these error bounds can be leveraged to verify the correctness of a model and its applicability to a given problem. Additionally, the direct dependence of these models and their bounds on the training data or simple basis functions, allows for predictions to be interpreted, particularly when working with small- to medium-sized data sets.
@@ 2 Interpolation Methods and Error Bounds
In this section we discuss common methods for interpolating an unknown blackbox function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ at an arbitrary “test” point given arbitrary training data. More precisely, we assume we have access to training data inputs ${\cal D}=\{x_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$ and paired outputs $f\left(\cal D\right)=\{y_{i}\}_{i=1}^{n}\subset\mathbb{R}$, with an arbitrary “test” or “query” point $q\in{\cal X}\subset\mathbb{R}^{d}$. We have chosen common methods from the literature that are well-defined for large values of $n$ and $d$, and do not rely upon any structured arrangement of ${\cal D}$. All of our analysis is also based upon scattered training data ${\cal D}$, which is the least restrictive case.
@@ 2.1.2 Advantages and challenges of Delaunay interpolants
A challenge when employing Delaunay interpolants is a potentially expensive computational cost at inference time. Many common approaches for computing Delaunay triangulations are impractical in even moderate dimensions, due to the exponential size of $DT({\cal D})$ with growing $d$ [35]. However, it is possible to identify the single simplex needed to calculate ${\hat{f}}_{DT}(q)$ for any given $q$ at inference time, for the cost of solving a linear program [36, 37]. This approach is scalable enough for usage with moderately-sized training sets, but still considerably more expensive than making a single prediction with many comparable neural network models, such as MLPs.
Additionally, when implementing such an approach, one must be careful of geometric degeneracy, which results in a degenerate linear program and could cause failures in certain solvers. These degeneracies occur with probability zero in theory for randomly scattered data, but are extremely common in real-world data sets. For instance, if any subset of a training set is grid-aligned or if all the data lies in a hyperplane of codimension at least 1, the geometry of the data would be considered degenerate. These degeneracies can be handled robustly. Grid-aligned data still has associated Delaunay interpolants, but there may be more than one valid choice. A hyperplane containing all the data can be detected and a Delaunay interpolant can be defined over this hyperplane.
@@ 2.2 RBF Interpolants and Thin-Plate Splines
The first step in fitting an RBF interpolant is (typically) fitting a “polynomial tail” $t(q)$, computed via generalized linear regression with polynomial basis functions. In this paper, we will use either a least-squares approximation for a linear tail, resulting in $t(q)=t_{1}(q):=\alpha^{\top}q+\beta$ for some $\alpha,\beta$, or a constant tail given by $t(q)=t_{0}(q):=\sum_{i=1}^{n}f(x^{(i)})/n$. It is possible to employ higher order techniques in some settings, such as RBF-FD approximations [39], but such methods are only feasible in low dimensions, typically $d\leq 3$.
@@ 2.2.1 Error bounds for RBFs
Error analysis for RBF interpolation is tricky because the error bounds depend on globally-dependent quantities derived from $f$, which would rarely be known in practice. We focus on a bound provided by Powell [41] and explained in [42, Theorem 1] as follows. Define
@@
Sticking with a mere Lipschitiz continuity assumption for $f$, we are not aware of a computable estimate for $C$ or $\|f\|_{\Phi}$. However, it is useful to note the scaling behavior of (10) with respect to $h$. Note that $h$ is a measure of the density of ${\cal D}$ in ${\cal X}$. If ${\cal D}$ becomes dense enough, the estimate is dominated by the $\log(h^{-1})$ term, which blows up as $h\rightarrow 0$, in that large $h$ values indicate less density, somewhere in $\mathcal{X}$. The blow up is a feature, not a bug, of the estimate, as it reflects the numerical ill-conditioning of the linear system (8) as the training data clusters.
@@ 2.2.2 Advantages and challenges of RBFs
When the size of the data set $n$ is small, RBF interpolants are relatively cheap to fit. Additionally, for many choices of basis functions—including the thin-plate spline functions that we will describe in the next section—there are often no hyperparameters. Evaluation of the interpolant at query-time is also quick with $O(nd)$ complexity. It is thus no surprise that RBF methods are a popular off-the-shelf solution to a wide variety of interpolation problems.
@@
Compared to the other interpolants that we consider, one critical challenge for a generic RBF interpolant is that the error bound cannot be easily calculated or approximated. One critical unknown in (10) is the value of the constants $C$ and $\|f\|_{\Phi}$. In principle, these values can be determined independently of the data set ${\cal D}$, however in practice they cannot be known for many real-world applications. We will see a special case in Section 2.3.1 where an error bound can be derived by statistical means, but this approach is kernel-dependent.
@@ 2.3.1 Approximate error bounds for GPs
In light of these assumptions, following the analysis in [16, Ch. 2], the posterior distribution for ${\hat{f}}_{GP}$ at $q$ is given by
$p\left({\hat{f}}_{GP}(q)~{}|~{}{\cal D}\right)\sim{\cal N}\left(\mathbb{E}\left[{\hat{f}}_{GP}\right](q),K(q,x)\right),$
@@
$K(q,x)=\tau_{f}\left(e^{-\|q-x\|_{2}^{2}/\tau}-\left(\phi^{(1)}(q),\ldots,\phi^{(n)}(q)\right)A_{GP}^{-1}\left[\begin{array}[]{c}\phi^{(1)}(x)\\ \vdots\\ \phi^{(n)}(x)\end{array}\right]\right),$
where $\tau_{f}$ is some estimate for the width of the range of $f$ and $A_{GP}$ is given by the RBF interpolation matrix from (8) with $\Phi=GP$. For an estimate of the variance at $q$, which can be used to evaluate the expected variance in $|f(q)-{\hat{f}}_{GP}(q)|$, this reduces to
$K(q,q)=\tau_{f}\left(1-\left\|\left(\phi^{(1)}(q),\ldots,\phi^{(n)}(q)\right)\right\|_{A^{-1}_{GP}}^{2}\right).$ (12)
Here, $\|\cdot\|_{A^{-1}_{GP}}$ is the 2-norm rescaled by $A^{-1}_{GP}$, which is also a norm since $A^{-1}_{GP}$ is symmetric positive definite.
Note that this interpretation of GPs relies on additional assumptions about the smoothness of $f$ that may not hold true in practice. For many real-world $f$, it is incorrect to assume that response values are Gaussian distributions and correlated with other response values based on the exponential of the inverse squared distance in the input space. It is for this reason that $K(q,q)$ is only an estimate for the actual variance. Further, if $f$ is not smooth enough, the error approximation by the above interpretation yields poor estimates. Nevertheless, the uncertainty estimate is widely used and is considered a useful heuristic in many applications [46].
@@ 2.3.3 Our methods for fitting GP interpolants
Instead, we solve a well-defined optimization problem based on the statistical literature to select a maximum likelihood estimate of $\tau$ via scikit-learn [48]. Specifically, the log-marginal likelihood for observing the response values $\{y_{i}\}_{i=1}^{n}$ given the data $\{x_{i}\}_{i=1}^{n}$ is
$\log p(y_{i}|x_{i})=-\frac{1}{2}\sum_{i=1}^{n}y_{i}c_{i}-\sum_{i=1}^{n}\log L_{ii}-\frac{n}{2}\log 2\pi~{}{=\frac{1}{2}\log\det\left(A_{GP}\right),}$
where $L$ is a Cholesky factor of $A_{GP}$ [49]. Recall that $c:=A^{-1}_{GP}f$ is a function of $\tau$ since $A_{GP}$ is a function of $\tau$. In scikit-learn, this marginal likelihood function is maximized over possible values of $\tau$ using a multi-start L-BFGS solver. While this is still technically a hyperparameter optimization problem, the problem is relatively inexpensive to solve and extremely reproducible since we are tuning a single variable using second-order gradient-based methods.
@@ 2.4 Artificial Neural Network and Multilayer Perceptron Models
where $u(\cdot)$ is a nonlinear activation function, $W^{(i)}$ is a tensor whose dimensions are compatible with $x^{(i)}$, and $b^{(i)}$ is a tensor whose dimension match $W^{(i)}x^{(i)}$.
@@ 3.1 Dataset generation and experimental setup
We have carried out numerous experiments exploring the interpolant errors and bounds over this parameter space and report the most relevant findings below. We found negligible variations by changing the spacing parameter and hence only report results using Sobol sequence sampling.
@@ 3.2 Effect of function variation ($\omega$) over $n$ and $d$
In Figure 2a, we show the results for fixing $n=8192$, with additional conventions as in the previous figure, for the Delaunay and MLP results. Notice that the relative positions of the error values are mostly stable for $d\geq 5$. We infer that the $d=5$ case from Figure 1a is representative of the relative performance of these interpolants for higher dimensions (at least up to 20). For $d<5$, we see smaller errors in the interpolation cases, due to the relative density of our fixed $n$ in low dimensions. As $d$ grows, the error effectively plateaus while the bound increases, reflecting the greater uncertainty inherent in the higher dimensional settings.
@@ 3.3 Effect of skew ($\alpha$) and the importance of scaling
3.4 Prevalence of geometric extrapolation regimes as $d$ increases
@@
In Figure 5, we show the results of the following experiment. For $n=2^{8+2i}$, $i=0,1,2,3$ (the same values used in Figures 13, and 4) we create a data set using a Sobol sequence of $n$ points in $[-1,1]^{d}$ for $d=2,5,8,11$. We set $\alpha=0$ (no skew). The value of $\omega$ is irrelevant as response values are not used in this study. For each $d$, we draw $2^{10}$ points uniformly from $[-1,1]^{d}$, which we call the “test points.” For each test point $q$, we determine whether $q$ lies inside or outside the convex hull of the data set and mark it as an interpolation or extrapolation query, respectively. We also compute $||q||$. Finally, we plot a histogram of the $||q||$ values, using 20 bins from 0 to $\sqrt{d}$ and color each bar according to the proportion of interpolation (blue) vs extrapolation (red) queries.
@@
Additionally, observe that as dimension increases, the distribution of $||q||$ values becomes more sharply clustered in the interior of the range of possible values (note $||q||\in\left[0,\sqrt{d}\right]$ for $q\in[-1,1]^{d}$). This phenomenon, known as concentration of volume, is sometimes overlooked in the data science community although it is well understood from the mathematical setting; see, for instance, [64, Theorem 3.2] and the references that follow. The concentration in $||q||$ values is due in part to the uniform sampling strategy we selected to draw the test points, however, we observed very similar results when using Latin hypercube or Sobol sequence sampling instead. Weighted sampling strategies could help distribute $||q||$ values more evenly, however, whether that is beneficial or advisable depends highly on the application context.
The classical geometric definition of “extrapolation” employed above is deterministic and dovetails naturally with the $L_{\infty}$-type bounds we have analyzed. In statistical learning, the term “out-of-distribution” addresses a similar notion, but a single data point can rarely be definitively declared to lie “outside” a distribution. Accordingly, statistical bounds are typically in averaged $L_{2}$ norms, which can mask geometric trends in the data.
@@ 4 Case Study: Predicting Lift-Drag Ratios from Airfoil Images
Our training data for this task is obtained from the UIUC Airfoil Coordinates Database [65]. This database provides a collection of coordinates for approximately 1,600 airfoils and their corresponding lift-to-drag ratios. These data correspond to a wide range of applications from low Reynolds number airfoils for UAVs and model aircraft to jet transports and wind turbines. Therefore, rapid predictions for the quantity of interest for this dataset corresponds to potential savings in predictive modeling across several different domains.
Original ImagesLatent Space $\mathbb{R}^{d}$ecoded ImagesLift/Drag RatioEncoderDecoderOriginal dataInterpolation methods
Figure 6: The UIUC Airfoil Coordinates Database provides a mapping from a collection of “Original Images” to a scalar “Lift/Drag Ratio.” We train an autoencoder (top row), then deploy interpolation methods as mappings from the “Encoded Images” to “Lift/Drag Ratio.” We use this framework to validate the choice of latent space for the autoencoder, assess the characteristics of a hold-out “test set” of images, and interpret results from the interpolation methods.
Attempting to predict a scalar value directly from image data is a regression rather than a classification task, suggesting an interpolation approach should be used. However, interpolating over the space of original images—$\mathbb{R}^{128\times 128}$—is well-beyond the feasible range of the interpolation methods considered here. Thus, we use a combined approach, as illustrated in Figure 6. First we train an autoencoder to embed the original images into a latent space $\mathbb{R}^{d}$ for some $d\in[2,4,6,8]$. Then, we deploy our interpolation methods to approximate the mapping from the latent space to a lift-drag ratio value, that is, from $\mathbb{R}^{d}$ to $\mathbb{R}$.
The autoencoder is implemented as follows. We utilize two-dimensional convolutional operations alternating with max-pooling operations to reduce the dimensionality of the input image successively via three hidden layers. Subsequently, the two-dimensional (but reduced) representation is flattened and acted on by fully-connected operations to reach the desired dimensionality in the latent space. This procedure is mirrored in the decoder where instead of MaxPooling operations, we utilize UpSampling operations to increase the spatial resolution of the image. In all our hidden layers, both in the convolutional layers as well as the fully-connected layers, we utilize the Swish activation function. The output layer is selected to have a linear activation. All our models are trained for 100 epochs with a learning rate of 0.001 using the ADAM optimizer, with an early stopping criterion that terminates optimization if there is no improvement in validation performance over 10 consecutive epochs. This criterion ensures robustness to overfitting. Additional details can be observed in our associated code [63].
The data set we use was constructed as follows. Using the $\sim$1600 airfoil coordinates, 6837 pairings of images and lift/drag ratios were collected (by changing the angle of attack) and shared in a publicly available repository [^3] https://github.com/ziliHarvey/CNN-for-Airfoil. From these parings, a set of 668 data points is held out and designated as the “test set.” From the remaining 6169 data points, we set aside 616 points as a “validation set,” then train an autoencoder on all or a subset of the 5,553 remaining non-test, non-validation points. Note that for all interpolation techniques and the ReLU MLP, we have used the same implementations described in Section 3 with input and output rescaling.
4.1 Validation of latent encoding dimension via interpolation methods
We now demonstrate how interpolation assessments in the framework of Figure 6 can help answer two related validation questions when training the autoencoder: “Was the choice of latent space dimension optimal?” and “Was a sufficient amount of training data used?” In this context, for an image-output pair $(x,y)$, we define:
[…]
For all the autoencoders we trained, the reconstruction error on the validation set was $\approx 10^{-4}$, meaning it provided no information on the quality of the representation learned for different latent dimensions. To gain such information, we compute the true error and the approximation error for the various interpolation types discussed. We will see that the true error surpasses the approximation error when the latent space dimension is too small and that both types of error inform whether adding more data to the training set improves the embedding.
Our experiments for these validation tasks are as follows. For each latent space dimension, we train an autoencoder with $p$ percent of the 5,553 training points for $p\in[10,20,\ldots,100]$. In each trained autoencoder and for each interpolation method and the ReLU MLP, we compute the true error and approximation error at each of the 616 points in the validation set. We then compute the relative mean absolute error (MAE) of each error type. We show the results for latent dimension 2, 4, 6, 8 in Figures 7, 8, 9, 10, respectively. In each figure, the left plots show MAE values as $p$ increases, at two different scales for the $y$-axis. The right scatterplots compare the true and approximation errors for the 616 validation points, with extrapolation points marked as X instead of dots.